Pressure is an expression of force exerted on a surface per unit area. The standard unit of pressure is the Pascal (Pa), inches of mercury (inHg) or bar.

(a) Compare between gauge pressure, absolute pressure and ambient pressure

(b) A rigid container of a volume 50m3 having an initial pressure of 100kPa and temperature of 250K. After gas is added into the container the final temperature is 400K. Find the final pressure inside the container?
1
Expert's answer
2020-12-14T12:18:03-0500

a) The difference between the two measurements is relatively easily clarified: in a gauge pressure measurement, it is always the difference from the current ambient pressure that is measured. However, this pressure changes with the weather and the height above sea level. An absolute pressure measurement measures the difference from the ideal or absolute vacuum. This is why this measurement is independent of environmental influences such as weather or altitude.

b)


"\frac{p_1}{T_1}=\frac{p_2}{T_2}\\\frac{100}{250}=\frac{p_2}{400}\\p_2=160\ kPa"